HOW DO YOU FIND A PARALLEL EQUATION FROM Y= 6X-5 THAT PASSES THROUGH THE POINT (-6,5)
Answer by fractalier(6550)

If the line you want is parallel to y = 6x - 5, it must have the same slope, namely 6.
Then apply the point-slope formula:
y - y1 = m(x - x1) where your (x1,y1) is (-6,5) and m=6...
y - 5 = 6(x -(-6))
y - 5 = 6(x + 6)
y - 5 = 6x + 36
y = 6x + 41